the length of a reactangle is 2 cm more than twice its width. The area of the rectangle is 60 cm^2 find the length and thw width of the rectangle

Answer by josgarithmetic(39617)

Same as your other one but just different value for area.
L=2w+2, wL=60

w(2w+2)=60
w(w+1)=30
w^2+w-30=0
(w-5)(w+6)=0